Decision Variables: B= # of Basic models to be produced, profit=$45/unit P= # of Premium models to be produced, profit=$120/unit Objective: MaxTotal Profit= Max $45 B + $120 P Constraints: Software Engineer (SE): 20 B + 5 P <= 2400 minutes QC Manager: 15 B + 30 P <= 2400 minutes Demand for B: B <= 100 units Demand for P: P <= 50 units.

Optimal production quantities are: a) B=100 & P=50 b) B=50 & P=60 c)B=60 & P=50 d) B=100 & P=0 e)none of these
Answer #1

Objective Function

Max Profit Z = 45B + 120P

Constraints

20B + 5P <= 2400

15B + 30P <= 2400

B <= 100

P <= 50

Let's solve using Graph -

The grey shaded region is the feasible area

Max can occur at extreme points A, B, C or D

A = [60, 50]
B = [100, 30]
C = [100, 0]
D = [0, 50]

ZA = 45*60 + 120*50 = 8700
ZB = 45*100 + 120*30 = 8100
ZC = 45*100 + 120*0 = 4500
ZD = 45*0 + 120*50 = 6000

Hence, max occurs at B = 60, P = 50

Hence, option (c) is correct option